# WildChat-50m: A Deep Dive Into the Role of Synthetic Data in Post-Training

### v1

@@ Abstract
Language model (LLM) post-training can refine behaviors and unlock new skills, but the open science supporting these post-training techniques is still in its infancy. One limiting factor has been the difficulty of conducting large-scale comparative analyses of synthetic data generating models and LLM judges. To close this gap, we introduce WildChat-50m, the largest public chat dataset to date. We extend the existing WildChat dataset to include responses not only from GPT, but from over 50 different open-weight models, ranging in size from 0.5B to 104B parameters. We conduct an extensive comparative analysis and demonstrate the potential of this dataset by creating Re-Wild, our own public SFT mix, which outperforms the recent Tulu-3 SFT mixture from Allen AI with only 40% as many samples. Our work is available at https://github.com/penfever/wildchat-50m.
@@ 1 Introduction
    We demonstrate the power of our dataset by using it as the basis of Re-Wild, a novel data mix for supervised fine-tuning (SFT) of LLMs. When we fine-tune Llama-3.1 8B Base on Re-Wild, wehow that our models outperform the SFT mix proposed in Tulu-3 (Lambert et al., 2024), along with several other existing, strong SFT baselines on a range of post-training benchmarks.
@@ 2.1 Data Collection
Although, to our knowledge, there are no large-scale diverse chat transcripts dataset of synthetically generated LLM responses, there are at least two recent large chat datasets available on which one could potentially base such a dataset: WildChat-1M from AllenAI, and LMSys-Chat-1M from LMSys (Zhao et al., 2024; Zheng et al., 2024). Although both have their strengths, we chose to focus on the former because of its rich variety of use-cases (including those which are potentially toxic), its diverse regional and temporal dimensions, and its relatively low levels of contamination for commonly used test sets (Zhao et al., 2024; Lambert et al., 2024).
@@ 2.2 Dataset Analysis
WildChat-50m currently includes 19 unique pre-trained models and 35 post-trained model variants, for a total of 54 checkpoints represented. With the exception of the responses in the original WildChat dataset, which are sourced from various GPT checkpoints, all responses in WildChat-50m are derived from models sourced from HuggingFace; the release dates range from July 2023 to November 2024, and the parameter counts range from 0.5B to 104B. For a comprehensive list of all the LLMs we used in the study, please refer to Sec. C. We attempted to select a diverse set of models; our main limiting factor was compatibility with our hardware setup, and with VLLM as an inference engine.
Naming Conventions. In order to make this paper more readable, we will employ certain naming conventions for the models we describe, enumerated below.
@@
It would be natural to assume, therefore, that LLMs that share neither pre-training nor post-training data would likewise may produce substantively different responses to prompts. However, our results show that this does not appear to be the case; LLM responses are unusually similar to one another. See Sec. A for a deeper analysis of this result, as well as every score (with associated standard deviation) for every model.
@@ 3.1 Re-Wild: A new data mixture for SFT
Training. We conduct our SFT experiments using a modified version of the Axolotl framework (Lian, 2025). We use the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of 2e-5, a single epoch, and a cosine learning rate scheduler, with eight steps of gradient accumulation, in bf16 precision. We also utilize several techniques to optimize training speed, such as gradient checkpointing, flash attention, and in some cases, FSDP (full shard, autowrap). Some artifacts (configurations, logs) are available on our GitHub repo, and others will be made available soon. Each of our SFT runs utilizes one 4xH100 node. The average time to fine-tune a model for 250,000 conversations takes approximately 5.5 hours.
@@
In order to make comparisons with past and future work easier, we select benchmarks which are popular and prominent in the recent research literature. We break down the concept of alignment into subcategories such as generalist chat capability, world knowledge, and instruction following. Following recent work, we select a mix of ground-truth benchmarks and LLM-judge benchmarks, in order to balance out the potential confounds inherent to each evaluation method (Feuer et al., 2024a; White et al., 2024). For generalist chat capabilities, we use MixEval, AlpacaEval2, and MTBench with GPT-4o-mini as the judge LLM (Ni et al., 2024; Dubois et al., 2024; Zheng et al., 2023). In some cases, we also report the average score over all of our benchmarks. For AlpacaEval2, we report length-controlled win rate. For instruction following and world knowledge, we utilize the recent HuggingFace OpenLLM Leaderboard 2 (Fourrier et al., 2024). For IFEval, we report prompt-level strict accuracy because it is more challenging and therefore better exhibits separation between models; however, we also include instance-level loose accuracy in our artifacts, where we observe generally similar trends. Finally, in some figures and tables, we report the average performance for all benchmarks (MixEval, AlpacaEval2-LC, MTBench / 10, OpenLLM LB 2) as Avg.
@@ 3.2 Key Findings
How much does the choice of data generating model impact downstream performance? Can we be certain that it is not the prompts, or perhaps some quirk of our training procedure, that have led to improved SFT performance using Re-Wild? To evaluate this concern, we compare the performance of six unique pretrained models from four distinct model families, including Qwen-2.5-72B-Instruct from Alibaba, Llama-3.3-70B-Instruct from Meta, Command-R-Plus from Cohere, and Jamba-1.5-Mini from AI21 (Qwen Team, 2024; Dubey et al., 2024; Durán et al., 2023; Lieber et al., 2024). The models range from 7B to 104B active parameters. The choice of DGM has a large effect, even when controlling for potential confounds such as the number of parameters for the model and the size of the context window. Furthermore, we find that no model dominates the benchmark, and that parameter count is not a perfect indicator of data quality. On three of the nine benchmarks we consider, the best performing model has fewer than 10B parameters. See Tab. 2.
@@
Do models benefit from blending DGMs? One impetus behind creating a very large prompt-response dataset like that of Zhao et al. (2024) is the intuition that more samples and more interactions will generally lead to better models. In the case of prompt diversity, this appears to be true (Feuer et al., 2024a), perhaps because it makes the model more robust to inconsistencies in prompting. But do the benefits of heterogeneous scaling extend to responses? In other words, do models trained on blended DGM responses outperform the sum of their parts? We conduct experiments to test this hypothesis, the results of which can be found in Sec. D, summarized here. We find that blending offers no benefit; the whole behaves almost exactly as the sum of its parts. This finding indicates that SDQ depends primarily on prompt diversity, and it is most effective to optimize, rather than generalize, responses.
@@
Is styling inherited during SFT? Effective use of presentational styling elements such as HTML tags, attributes and inline properties can have a strong effect on a text’s readability and clarity. We investigate whether such formal styling behaviors are inherited from the DGM during the SFT process. We select a subset of 80 turns from MTBench and examine the behavior of two DGMs (Qwen 2.5 72B and Llama 3.3 70B) and their finetunes. We convert the model responses from Markdown (which they commonly use in their responses) to HTML and report the absolute and proportional frequency of each styling tag. Absolute frequency is the raw count for each feature (see rows 1:4). Proportional frequency, for rows $A$, $B$, and feature $F$, is given by $A[F]\div B[F]$. The closer this quantity is to 1, the more similar the model responses.
@@
Do models learn better from DGMs in the same model family? Recent work from Tajwar et al. (2024) has shown that approaches that use on-policy sampling or attempt to push down the likelihood on certain responses (i.e., employ a “negative gradient”) outperform offline and maximum likelihood objectives. In this section, we inquire whether this finding extends not only to direct on-policy sampling, but something we might call approximate on-policy sampling, where a pretrained base model is fine-tuned on its own post-trained outputs, or those of a model from a similar family. Indeed, when we experiment with Qwen-2-7B and Llama-3-8B, we find that this approach produces stronger benchmark results when controlling for dataset size, and appears also to extend to a larger post-trained checkpoint in the same model family; see Sec. B for the complete results.
@@ 6 Conclusions
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work. In particular, we wish to acknowledge the fact that we are building upon a dataset (WildChat) which is known to contain examples of user inputs on potentially upsetting topics, including but not limited to sex, violence, and bigoted claims (Zhao et al., 2024). Following the authors of that work, we will require approval before researchers are allowed to access our datasets, and will release our data under the AI2 ImpACT License, which explicitly forbids certain use cases for this data.
In addition, our data will be subject to any additional restrictions from the licenses of the models we utilize. We also preserve, without modification, the original user inputs from WildChat (Zhao et al., 2024), and therefore in our work there remains the possibility that users may have inadvertently included personal information within their conversations which was not detected by the (considerable) safeguards put in place by the WildChat authors. We will make this clear to users upon dataset release.
@@ Appendix C List of All LLMs in the Study, with Citations
  • c4ai-command-r-plus-08-2024, from Durán et al. (2023)

### v2

@@ Abstract
Language model (LLM) post-training, from DPO to distillation, can refine behaviors and unlock new skills, but the open science supporting these post-training techniques is still in its infancy. One limiting factor has been the difficulty of conducting large-scale comparative analyses of synthetic data generating models and LLM judges. To close this gap, we introduce WildChat-50m, the largest public chat dataset to date. We extend the existing WildChat dataset to include responses not only from GPT, but from over 50 different open-weight models, ranging in size from 0.5B to 104B parameters. We conduct an extensive comparative analysis and demonstrate the potential of this dataset by creating Re-Wild, our own public SFT mix, which outperforms the recent Tulu-3 SFT mixture from Allen AI with only 40% as many samples. Our dataset, samples and code are available at https://github.com/penfever/wildchat-50m.
Machine Learning, ICML
@@ 1 Introduction
    We demonstrate the power of our dataset by using it as the basis of Re-Wild, a novel data mix for supervised fine-tuning (SFT) of LLMs. When we fine-tune Llama-3.1 8B Base on Re-Wild, we show that our models outperform the SFT mix proposed in Tulu-3 (Lambert et al., 2024), along with several other existing, strong SFT baselines on a range of post-training benchmarks.
@@ 2.1 Data Collection
Although, to our knowledge, there are no large-scale diverse chat transcripts dataset of synthetically generated LLM responses, there are at least two recent large chat datasets available on which one could potentially base such a dataset: WildChat-1M from AllenAI, and LMSys-Chat-1M from LMSys (Zhao et al., 2024b; Zheng et al., 2024). Although both have their strengths, we chose to focus on the former because of its rich variety of use-cases (including those which are potentially toxic), its diverse regional and temporal dimensions, and its relatively low levels of contamination for commonly used test sets (Zhao et al., 2024b; Lambert et al., 2024).
@@ 2.2 Dataset Analysis
WildChat-50m collects data from 19 unique pre-trained models (each of which is post-trained) and 35 post-trained model variants (with non-unique pre-trained models). This yields a total of 54 DGMs represented. With the exception of the responses in the original WildChat dataset, which are sourced from various GPT checkpoints, all responses in WildChat-50m are derived from models sourced from HuggingFace; the release dates range from July 2023 to November 2024, and the parameter counts range from 0.5B to 104B. For a comprehensive list of all the LLMs we used in the study, please refer to Sec. C. We attempted to select a diverse set of models; our main limiting factor was compatibility with our hardware setup, and with VLLM as an inference engine. The resulting public artifact is more than 50 times larger than the next largest public chat datasets of which we are aware, WildChat-1M and LMSys-Chat-1M.
Naming Conventions. In order to make this paper more readable, we will employ certain naming conventions for the models and datasets generated using DGMs described in this paper. The aforementioned conventions are enumerated below.
@@
It would be natural to assume, therefore, that LLMs that do not entirely share either pre-training or post-training data would likewise produce substantively different responses to prompts. However, our results show that this does not appear to be the case; LLM responses are unusually similar to one another. See Sec. A for a deeper analysis of this result, as well as every score (with associated standard deviation) for every model.
@@ 3 SFT Experiments
Our dataset, samples and code are available at https://github.com/penfever/wildchat-50m.
@@ 3.1 Re-Wild: A new data mixture for SFT
Training. We conduct our SFT experiments using a modified version of the Axolotl framework (Lian, 2025). We use the AdamW optimizer (Loshchilov & Hutter, 2017) with a learning rate of 2e-5, a single epoch, and a cosine learning rate scheduler, with eight steps of gradient accumulation, in bf16 precision. We also utilize several techniques to optimize training speed, such as gradient checkpointing, flash attention, and in some cases, FSDP (full shard, autowrap). The base model trained is always llama 3.1 8B (for us and baselines). All artifacts are available on our GitHub repo. Each of our SFT runs utilizes one 4xH100 node. The average time to fine-tune a model for 250,000 conversations takes approximately 5.5 hours.
@@
In order to make comparisons with past and future work easier, we select benchmarks which are popular and prominent in the recent research literature. We break down the concept of alignment into subcategories such as generalist chat capability, world knowledge, and instruction following. Following recent work, we select a mix of ground-truth benchmarks and LLM-judge benchmarks, in order to balance out the potential confounds inherent to each evaluation method (Feuer et al., 2024a; White et al., 2024). For generalist chat capabilities, we use MixEval, AlpacaEval2, and MTBench with GPT-4o-mini as the judge LLM (Ni et al., 2024; Dubois et al., 2024; Zheng et al., 2023). In some cases, we also report the average score over all of our benchmarks. For AlpacaEval2, we report length-controlled win rate. For instruction following and world knowledge, we utilize the recent HuggingFace OpenLLM Leaderboard 2 (Fourrier et al., 2024). For IFEval, we report prompt-level strict accuracy because it is more challenging and therefore better exhibits separation between models; however, we also include instance-level loose accuracy in our artifacts, where we observe generally similar trends. Finally, in some figures and tables, we report the average performance for all benchmarks (MixEval, AlpacaEval2-LC, MTBench / 10, OpenLLM LB 2) as Avg. Where possible, we include 95% confidence intervals using the normal approximation method.
@@ 3.2 Key Findings
How much does the choice of data generating model impact downstream performance? Can we be certain that it is not the prompts, or perhaps some quirk of our training procedure, that have led to improved SFT performance using Re-Wild? To evaluate this concern, we compare the performance of six unique pretrained models from four distinct model families, including Qwen-2.5-72B-Instruct from Alibaba, Llama-3.3-70B-Instruct from Meta, Command-R-Plus from Cohere, and Jamba-1.5-Mini from AI21 (Qwen Team, 2024; Dubey et al., 2024; Gomez, ; Lieber et al., 2024). The models range from 7B to 104B active parameters. The choice of DGM has a large effect, even when controlling for potential confounds such as the number of parameters for the model and the size of the context window. Furthermore, we find that no model dominates the benchmark, and that parameter count is not a perfect indicator of data quality. On three of the nine benchmarks we consider, the best performing model has fewer than 10B parameters. See Tab. 2.
@@
Do models benefit from blending DGMs? One impetus behind creating a very large prompt-response dataset like that of Zhao et al. (2024b) is the intuition that more samples and more interactions will generally lead to better models. In the case of prompt diversity, this appears to be true (Feuer et al., 2024a), perhaps because it makes the model more robust to inconsistencies in prompting. But do the benefits of heterogeneous scaling extend to responses? In other words, do models trained on blended DGM responses outperform the sum of their parts? We conduct experiments to test this hypothesis, the results of which can be found in Sec. D, summarized here. We find that blending offers no benefit; the whole behaves almost exactly as the sum of its parts. This finding indicates that the dependencies on prompt diversity discovered in prior work do not extend to the space of model responses even for a large set of benchmarks. It is therefore most effective to optimize, rather than generalize, responses.
@@
Is styling inherited during SFT? Effective use of presentational styling elements such as HTML tags, attributes and inline properties can have a strong effect on a text’s readability and clarity. We investigate whether such formal styling behaviors are inherited from the DGM during the SFT process. We select a subset of 80 turns from MTBench and examine the behavior of two DGMs (Qwen 2.5 72B and Llama 3.3 70B) and their finetunes. We convert the model responses from Markdown (which they commonly use in their responses) to HTML and report the absolute and proportional frequency of each styling tag. Absolute frequency is the raw count for each feature (see Tab. 3 rows 1:4). Proportional frequency, for rows $A$, $B$, and feature $F$, is given by $A[F]\div B[F]$. The closer this quantity is to 1, the more similar the model responses.
@@
Do models learn better from DGMs in the same model family? Recent work from Tajwar et al. (2024) has shown that approaches that use on-policy sampling or attempt to push down the likelihood on certain responses (i.e., employ a “negative gradient”) outperform offline and maximum likelihood objectives. In this section, we inquire whether this finding extends not only to direct on-policy sampling, but something we might call approximate on-policy sampling, where a pretrained base model is fine-tuned on its own post-trained outputs, or those of a model from a similar family. Indeed, when we experiment with Qwen-2-7B and Llama-3-8B, we find that this approach produces stronger benchmark results when controlling for dataset size, and appears also to extend to a larger post-trained checkpoint in the same model family; see Sec. B for the complete results. We acknowledge that our experiments here are limited in scope, and consider this an important area of future study.
@@ 4 Related Work
Recently, Zhao et al. (2024a) introduced WildChat-1M, a dataset containing 1 million ChatGPT interaction logs collected in real-world settings. Their work takes a different approach to ours by focusing on capturing authentic human-AI interactions rather than generating synthetic conversations. WildChat-1M leverages multiple GPT model variants, with 76% of conversations using the GPT-3.5-Turbo API (including versions 3.5-turbo-0613, 3.5-turbo-0301, and 3.5-turbo-0125) and 24% utilizing the GPT-4 API (including 4-1106-preview, 4-0314, and 4-0125-preview). For human prompt acquisition, they deployed chatbot services on Hugging Face Spaces and implemented a two-step consent process, collecting anonymized data from users who explicitly opted into the research. While WildChat-1M offers valuable insights into real-world chatbot interactions across multiple languages, our approach with WildChat-50m explores the complementary direction of high-volume synthetic data for post-training.
@@ 6 Conclusions
This paper presents work whose goal is to advance the field of machine learning. There are many potential societal consequences of our work. In particular, we wish to acknowledge the fact that we are building upon a dataset (WildChat) which is known to contain examples of user inputs on potentially upsetting topics, including but not limited to sex, violence, and bigoted claims (Zhao et al., 2024b). Following the authors of that work, we will require approval before researchers are allowed to access our datasets, and will release our data under the AI2 ImpACT License, which explicitly forbids certain use cases for this data.
In addition, our data will be subject to any additional restrictions from the licenses of the models we utilize. We also preserve, without modification, the original user inputs from WildChat (Zhao et al., 2024b), and therefore in our work there remains the possibility that users may have inadvertently included personal information within their conversations which was not detected by the (considerable) safeguards put in place by the WildChat authors. We will make this clear to users upon dataset release.
@@ Appendix C List of All LLMs in the Study, with Citations
  • c4ai-command-r-plus-08-2024, from Gomez
@@ Appendix E Output Samples
Appendix F Common Pitfalls and Best Practices in Synthetic Data Generation
Synthetic data generation for large language model training presents numerous challenges that can significantly impact downstream performance. Through our experience developing WildChat-50m, we identified several critical pitfalls that researchers should avoid when creating large-scale synthetic datasets.
One of the most consequential errors is bias transfer from data-generating models (DGMs). When a DGM has been trained on biased data or exhibits systemic failures in specific domains, these issues propagate into the synthetic dataset and may be further amplified when models are trained on this data. In our work, we observed that using a diverse mix of DGMs helps mitigate this problem by ensuring no single model’s biases dominate the dataset. Quantitative bias evaluation should be performed on sample generations before committing to full-scale data generation, as discovering bias issues after generating millions of examples can necessitate discarding entire dataset subsets—a tremendously wasteful outcome in terms of both computational resources and research timelines.
Technical configuration issues can equally undermine synthetic data quality. Inadequate context window sizing, improper temperature settings, poor prompt engineering, and insufficient runtime validation can all produce fatally flawed datasets. We found that temperature settings substantially influence output diversity and quality—temperatures that are too low (¡ 0.3) lead to repetitive, generic responses while settings that are too high (¿ 1.2) frequently introduce hallucinations, grammatical errors, and incoherence. During our data generation, we witnessed how improper prompt formatting with missing or inconsistent system instructions caused several models to produce unusable outputs, requiring us to restart entire batches. Another critical technical consideration is tokenization differences between models; in our collection process, we observed that using identical maximum token limits across architecturally diverse models led to dramatically different response lengths, necessitating model-specific calibration.
Based on our experience, we recommend the following best practices: (1) Conduct small-scale pilot runs with diverse inputs before full production to identify potential issues; (2) Implement comprehensive runtime validation including token count verification, response coherence checks, and completion confirmation; (3) Meticulously document all hyperparameters used in generation across different model architectures; (4) Employ robust error handling and logging systems that can gracefully recover from inference failures without losing progress; and (5) Include a diverse array of DGMs to prevent any single model’s limitations from dominating the resulting dataset. Following these practices can substantially improve synthetic data quality while avoiding costly regeneration cycles that consume valuable computational resources.